# Asynchronous Large Language Model Enhanced Planner for Autonomous Driving

### v2

@@ Abstract
Despite real-time planners exhibiting remarkable performance in autonomous driving, the growing exploration of Large Language Models (LLMs) has opened avenues for enhancing the interpretability and controllability of motion planning. Nevertheless, LLM-based planners continue to encounter significant challenges, including elevated resource consumption and extended inference times, which pose substantial obstacles to practical deployment. In light of these challenges, we introduce AsyncDriver, a new asynchronous LLM-enhanced closed-loop framework designed to leverage scene-associated instruction features produced by LLM to guide real-time planners in making precise and controllable trajectory predictions. On one hand, our method highlights the prowess of LLMs in comprehending and reasoning with vectorized scene data and a series of routing instructions, demonstrating its effective assistance to real-time planners. On the other hand, the proposed framework decouples the inference processes of the LLM and real-time planners. By capitalizing on the asynchronous nature of their inference frequencies, our approach have successfully reduced the computational cost introduced by LLM, while maintaining comparable performance. Experiments show that our approach achieves superior closed-loop evaluation performance on nuPlan’s challenging scenarios.
@@ 1 Introduction
As illustrated in Fig.1(a), current learning-based real-time motion planning frameworks Kendall et al. (2019); Li et al. (2023); Hallgarten et al. (2023); Renz et al. (2022); Scheel et al. (2022); Huang et al. (2023) typically utilize vectorized map information as input and employ a decoder to predict trajectories. As purely data-driven methods, they are particularly vulnerable to long-tail phenomena, where their performance can significantly degrade in rare or unseen scenarios Chen et al. (2022). Moreover, while some rule-based strategies exist, their manual crafting of rules are found to be inadequate for capturing the entirety of potential complex scenarios, resulting in driving strategies that tend towards extremes either excessive caution or aggression. Furthermore, both learning-based and rule-based planning frameworks suffer from low controllability, which raises concerns regarding the safety and reliability of such systems in dynamic environments.
Recently, the considerable potential of Large Language Models (LLMs), including GPT-4 Achiam et al. (2023) and Llama2 Touvron et al. (2023), has been extensively explored within the realm of autonomous driving. Their extensive pre-training on large-scale datasets has established a robust foundation for comprehending traffic rules and scenarios. Consequently, LLM-based planners have demonstrated superior performance in scene analysis, reasoning, and human interaction, heralding new prospects for enhancing the interpretability and controllability of motion planning Zhou et al. (2023); Yang et al. (2023); Cui et al. (2024). Nonetheless, as shown in Fig.1(b), these models frequently encounter several of these specific challenges: 1) The scene information is described through language, which could be constrained by the permissible input token length, making it challenging to encapsulate complex scene details comprehensively and accurately Mao et al. (2023a); Sha et al. (2023); Wen et al. (2023); Mao et al. (2023b). 2) Prediction via linguistic outputs entails either directly issuing high-level commands that are then translated into control signals, potentially leading to inaccuracies, or outputting trajectory points as floating-point numbers through language, a task at which LLMs are not adept Xu et al. (2023); Mao et al. (2023b); Keysan et al. (2023). 3) Prevalent frameworks primarily utilize LLMs as the core decision-making entity. While this strategy offers advantages in performance, the inherently large number of parameters in LLMs results in a noticeable decrease in inference speed relative to real-time planners, presenting substantial obstacles to their real-world implementation.
@@ 2.1 Motion Planning For Autonomous Driving
The classic modular pipeline for autonomous driving contains perception, prediction, and planning. Within this framework, the planning stage leverages the outputs of perception to determine the autonomous vehicle’s behavior and formulate a future trajectory, which is subsequently executed by the control system. The modular pipeline architecture has been extensively adopted across the industry, notably in frameworks like Apollo Baidu (2019). Distinct from end-to-end approaches Hu et al. (2023, 2022), modular architecture facilitates research concentration on individual tasks by delineating the data interfaces among modules.
Autonomous driving planners can be mainly divided into rule-based and learning-based planners. Rule-based planners Treiber et al. (2002); Kesting et al. (2007); Dauner et al. (2023); Fan et al. (2018); Urmson et al. (2008); Leonard et al. (2008); Bacha et al. (2008) operate on a set of predefined rules to determine the vehicle’s future trajectory, including protocols for maintaining a safe following distance and adherence to traffic signals. IDM Treiber et al. (2002) ensures that the ego car maintains a minimum safe distance from the leading vehicle. It incorporates a formula to calculate an appropriate speed that factors in the braking distance of the preceding vehicle, the stopping distance of the following vehicle, and the necessary safety margin. PDM Dauner et al. (2023) extends IDM by selecting the IDM proposal with the highest score as the final trajectory. This model has achieved state-of-the-art performance in the closed-loop evaluation of the nuPlan Challenge 2023 Caesar et al. (2021). Nonetheless, rule-based planners often struggle to effectively manage complex driving scenarios that fall outside the scope of their predefined rule sets.
Learning-based planners Kendall et al. (2019); Li et al. (2023); Hallgarten et al. (2023); Renz et al. (2022); Scheel et al. (2022); Huang et al. (2023) try to mimic human expert driving trajectories from large-scale driving datasets through imitation learning or offline reinforcement learning. Nonetheless, limited by the breadth of the datasets and the complexity of the models, these planners still have significant room for improvement in aspects such as routing information comprehension, environmental awareness, and decision-making.
@@ 2.2 LLM For Autonomous Driving
The rapid advancement of Large Language Models has been noteworthy. These models, including GPT-4 Achiam et al. (2023) and Llama2 Touvron et al. (2023), have been trained on extensive textual datasets and exhibit exceptional generalization and reasoning abilities. A growing body of research has explored the application of LLMs’ decision-making capacities to the domain of autonomous driving planning Sima et al. (2023); Chen et al. (2023b); Mao et al. (2023a); Sha et al. (2023); Jin et al. (2023); Wen et al. (2023); Mao et al. (2023b); Wang et al. (2023b); Shao et al. (2023); Sharan et al. (2023); Han et al. (2024); Xu et al. (2023); Fu et al. (2024); Cui et al. (2023); Wang et al. (2023a); Ma et al. (2023a); Nie et al. (2023); Ma et al. (2023b); Chen et al. (2023a); Liu et al. (2023).
Some efforts Fu et al. (2024); Mao et al. (2023b); Wen et al. (2023); Mao et al. (2023a) have explored the integration of scene information—encompassing the ego vehicle’s status, as well as information of obstacles, pedestrians, and other vehicles—into LLMs using linguistic modalities to facilitate decision-making and generate both decision and explanations. These approaches encounter limitations due to finite context length, challenging the encoding of sufficiently precise information for effective decision-making and reasoning. To address these constraints, multi-modal strategies, exemplified by DrivingWithLLM Chen et al. (2023b) and DriveGPT4 Xu et al. (2023), have been developed. These methods align vectorized or image/video modalities with linguistic instructions, enabling a more comprehensive interpretation of driving scenarios. However, these methods utilize language to express control instructions, which have certain limitations. DrivingWithLLM outputs high-level commands in linguistic expressions, which, despite enhancing question-and-answer (QA) interaction performance, compromises the fidelity of translating complex reasoning into precise vehicle control. DriveGPT4 articulates waypoints through language, demonstrating commendable open-loop performance, yet it lacks evaluation in closed-loop simulation.
Furthermore, some efforts Sha et al. (2023); Shao et al. (2023) are aimed at closed-loop evaluation, by connecting low-level controllers or regressors behind the large language model to achieve precise vehicular control. LMDrive Shao et al. (2023) employs a sequence of continuous image frames, supplemented by navigation instructions, to realize closed-loop driving capabilities, but still requires a complete process of LLM inference in each planning step. LanguageMPC Sha et al. (2023) uses LLM to obtain Model Predictive Control parameters to achieve control in a training-free manner. Additionally, these methods invariably require serial decoding of language at every planning step, or the very least, a complete inference of LLM, which presents challenges in real-time responsiveness and limits their practical deployment and application.
In our approach, we shift the focus from utilizing LLM for direct language output to scaling up real-time learning-based planners. This strategy aims to enhance the existing learning-based planners’ environmental comprehension. Moreover, our framework allows for the LLM and real-time planners to operate in a decoupled manner, executing planning at distinct inference rates. This strategy effectively mitigates the LLMs’ inference latency, thereby smoothing the path towards real-world deployment.
@@ 3 Data Generation
The nuPlan Caesar et al. (2021) dataset is the first large-scale benchmark for autonomous driving planning, containing a total of $1,200$ hours of human driving data in real-world in four cities: Boston, Pittsburgh, Las Vegas, and Singapore. To accommodate different stages of training, we developed pre-training and fine-tuning datasets derived from nuPlan Train set and Val set. Specifically, we selected only the data corresponding to $14$ official challenging Motional (2023) scenario types.
@@ 3.1 Pre-training Data Generation
To augment LLM’s comprehension of instructions in the context of autonomous driving, we designed a dataset comprising solely of language-based QAs, which is aligned with LLM’s inherent modality to facilitate more effective assimilation of instruction semantics. The pre-training dataset includes two components: Planning-QA and Reasoning1K. Samples of both datasets are shown in the supplementary material.
@@ 3.1.1 Planning-QA
is crafted using a rule-based approach, thus facilitating the scalability of the dataset. It is strategically designed to augment the LLM grasp of the interrelations among waypoint - high-level instructions - control. Within this framework, the waypoint constitutes an array of points, whereas high-level instructions constructed by velocity commands (encompassing stop, accelerate, decelerate, maintain speed) and routing commands (including turn left, turn right, go straight), the aspect of control predominantly pertains to assess the values of velocity and acceleration. Planning-QA encapsulates $6$ types of questions, each of which focuses on the conversion between waypoint - high-level instructions - control.
@@ 3.2 Fine-tuning Data Generation
Aiming to achieve multi-modal understanding and alignment further, we constructed fine-tuning data based on $100,000$ scenarios, capturing one frame every $8$ seconds, to compile a training set of $180,000$ frames and a validation set of $20,000$ frames, with each sample incorporating both map data and linguistic prompt to construct multi-modal information. Crucially, the scenario type distribution within both training and validation datasets adheres to the distribution of the entire nuPlan trainval dataset.
For the extracted vectorized scene information, similar to Huang et al. (2023), ego information and $20$ surrounding agent information over $20$ historical frames, in addition to global map data centered around the ego are involved. For agents that cannot obtain complete $T$ historical features, missing information is supplemented through interpolation. Regarding map data, three categories of geometric information within a predetermined radius $R=60$ are extracted: lane center lines, crosswalks, and road lanes.
LLM’s prompt is comprised of two parts: system prompt and series of routing instructions. In the template, we propose <map> serves as a placeholder for map data, intended to be replaced with vectorized scene information upon input. Concerning routing instructions, a rule-based approach is employed to transform the pathway into a series of instructions enhanced with distance information, which encompasses four types of actions: go straight, turn right, turn left, and stop. Notably, the stop action would not be utilized to denote cessation activities but is exclusively reserved for scenarios where the entire path is less than $0.5$ meters. Regarding training dataset preparation, the ego vehicle’s ground truth trajectory over the ensuing $8$ seconds is harnessed as the pathway for routing instruction generation. During simulation, based on the observation of the current scene, a pathway that adheres to a specified maximum length is found through a hand-crafted method as a reference path for instruction generation.
@@ 4 Methodology
In this work, we introduce the asynchronous LLM-enhanced closed-loop framework, AsyncDriver. As illustrated in Fig. 2, our approach mainly includes two components: 1) The Scene-Associated Instruction Feature Extraction Module, where multi-modal inputs, comprising vectorized scene information and series of routing instructions, are aligned through the Alignment Assistance Module, facilitating the extraction of high-level instruction features; 2) The Adaptive Injection Block, which receives the former feature as a guide for the real-time planner to get more precise and controllable trajectory prediction. Additionally, owning to the nature of our framework, the inference frequency between the LLM and the real-time planner could be decoupled and regulating by asynchronous interval, which markedly enhancing inference efficiency.
@@ 4.1.1 Multi-modal Input
In each planning iteration, the vectorized scene information is acquired from the simulation environment. Analogous to the approach employed by GameFormer Huang et al. (2023), historical trajectory and state information of both the ego and other agents are extracted, alongside global map data. Vectorized scene information for the real-time planner is provided in the same manner. All vector data are relative to the position of the ego. Subsequently, through the processing by the Vector Map Encoder and Map Adapter, we derive map embeddings. These map embeddings, along with language embeddings, are then fed into the Llama2-13B backbone to obtain the final hidden features $h=\{h_{0},h_{1},...,h_{-1}\}\in\mathbb{R}^{N_{h}\times D_{llm}}$.
@@ 4.2 Adaptive Injection Block
We adopt the decoder structure of Huang et al. (2023) as our basic decoder layer, facilitating the adaptive integration of scene-associated instruction features by evolving the conventional transformer-based decoder layer into an Adaptive Injection Block.
@@ 4.4 Training Details
Following Huang et al. (2023), the planning loss comprises two parts: 1) Mode prediction loss. $m$ different modes of trajectories of neighbor agents are represented by Gaussian mixture model (GMM). For each mode, at any given timestamp $t$, its characteristics are delineated by a mean ${\mu}^{t}$ and covariance ${{\sigma}^{t}}$ forming a Gaussian distribution. The best mode ${m}^{*}$ is identified through alignment with the ground truth and refined via the minimization of negative log-likelihood. 2) Ego Trajectory Prediction loss. Future trajectory points of ego vehicle are predicted and are refined by $l1$ loss. Consequently, the planning loss is as follows:
@@ 5.1.1 Evaluation Settings
In accordance with the settings of nuPlan challenge 2023 Motional (2023), we select $14$ official challenging scenario types for training and evaluation. Nevertheless, nuPlan Caesar et al. (2021) encompasses a vast collection of $757,844$ scenarios, considering that on one hand, the majority of simple scenarios fail to effectively assess the critical performance of planners, and on the other hand, the sheer volume of data leads to prolonged evaluation times, we have meticulously selected the Hard20 dataset. Concretely, we randomly picked $100$ scenarios for each type from the test set and employing the PDM Dauner et al. (2023) planner, which is the champion of nuPlan 2023 challenge, and preserved $20$ scenarios with the lowest scores per type, forming Hard20 as our test set consisting of a total of $279$ scenarios.
@@
UrbanDriver Scheel et al. (2022) 35.35 75.53 97.12 98.56 85.23 55.21 81.62 47.84
GCPGP Hallgarten et al. (2023) 36.85 81.29 98.20 77.33 46.96 72.30 97.92 68.34
IDM Treiber et al. (2002) 53.07 84.94 98.02 83.15 64.79 74.01 96.38 60.57
GameFormer Huang et al. (2023) 62.05 93.54 98.74 83.15 66.27 86.02 98.19 74.55
PDM-Hybird Dauner et al. (2023) 64.05 95.34 99.10 75.98 67.93 87.81 99.57 72.75
PDM-Closed Dauner et al. (2023) 64.18 95.69 99.10 77.06 68.20 87.81 99.57 73.47
@@
Table 2: Scores per scenario types in evaluation on nuPlan Closed-Loop Reactive Challenges on Hard20 split. The best results are highlighted in bold , while the second-best results are underscored with an underline for clear distinction. Types 00-$13$ represent the $14$ official scenario types of the nuPlan challenge 2023 Motional (2023), with specific details provided in the supplementary materials.
@@
UrbanDriver Scheel et al. (2022) 69.39 15.78 44.59 7.42 13.7 27.14 0.00 19.44 20.8 23.61 68.33 93.16 33.78 56.39
GCPGP Hallgarten et al. (2023) 59.50 35.91 33.6 29.33 42.84 17.24 4.86 32.33 8.22 36.23 71.32 80.46 36.19 23.61
IDM Treiber et al. (2002) 70.69 44.84 91.54 54.08 50.22 41.71 4.76 53.97 37.53 60.33 83.97 93.03 45.77 12.44
GameFormer Huang et al. (2023) 84.32 65.78 83.62 49.03 71.79 36.8 0.00 51.76 55.03 77.24 82.83 98.24 49.41 56.16
PDM-Hybird Dauner et al. (2023) 87.68 69.61 87.20 49.95 82.80 41.32 4.23 54.98 51.72 82.95 80.37 98.40 37.77 71.99
PDM-Closed Dauner et al. (2023) 87.67 70.93 87.20 49.95 82.80 41.25 4.22 53.53 51.57 82.96 80.37 98.40 39.95 72.04

### v3

@@ Abstract
Despite real-time planners exhibiting remarkable performance in autonomous driving, the growing exploration of Large Language Models (LLMs) has opened avenues for enhancing the interpretability and controllability of motion planning. Nevertheless, LLM-based planners continue to encounter significant challenges, including elevated resource consumption and extended inference times, which pose substantial obstacles to practical deployment. In light of these challenges, we introduce AsyncDriver, a new asynchronous LLM-enhanced closed-loop framework designed to leverage scene-associated instruction features produced by LLM to guide real-time planners in making precise and controllable trajectory predictions. On one hand, our method highlights the prowess of LLMs in comprehending and reasoning with vectorized scene data and a series of routing instructions, demonstrating its effective assistance to real-time planners. On the other hand, the proposed framework decouples the inference processes of the LLM and real-time planners. By capitalizing on the asynchronous nature of their inference frequencies, our approach have successfully reduced the computational cost introduced by LLM, while maintaining comparable performance. Experiments show that our approach achieves superior closed-loop evaluation performance on nuPlan’s challenging scenarios. The code and dataset are available at https://github.com/memberRE/AsyncDriver.
Keywords:
Autonomous Driving Large Language Models Motion Planning
@@ 1 Introduction
As illustrated in Fig.1(a), current learning-based real-time motion planning frameworks [20, 24, 13, 32, 33, 18] typically utilize vectorized map information as input and employ a decoder to predict trajectories. As purely data-driven methods, they are particularly vulnerable to long-tail phenomena, where their performance can significantly degrade in rare or unseen scenarios [6]. Moreover, while some rule-based strategies exist, their manual crafting of rules are found to be inadequate for capturing the entirety of potential complex scenarios, resulting in driving strategies that tend towards extremes either excessive caution or aggression. Furthermore, both learning-based and rule-based planning frameworks suffer from low controllability, which raises concerns regarding the safety and reliability of such systems in dynamic environments.
Recently, the considerable potential of Large Language Models (LLMs), including GPT-4 [1] and Llama2 [38], has been extensively explored within the realm of autonomous driving. Their extensive pre-training on large-scale datasets has established a robust foundation for comprehending traffic rules and scenarios. Consequently, LLM-based planners have demonstrated superior performance in scene analysis, reasoning, and human interaction, heralding new prospects for enhancing the interpretability and controllability of motion planning [48, 46, 9]. Nonetheless, as shown in Fig.1(b), these models frequently encounter several of these specific challenges: 1) The scene information is described through language, which could be constrained by the permissible input token length, making it challenging to encapsulate complex scene details comprehensively and accurately [28, 34, 44, 29]. 2) Prediction via linguistic outputs entails either directly issuing high-level commands that are then translated into control signals, potentially leading to inaccuracies, or outputting trajectory points as floating-point numbers through language, a task at which LLMs are not adept [45, 29, 22]. 3) Prevalent frameworks primarily utilize LLMs as the core decision-making entity. While this strategy offers advantages in performance, the inherently large number of parameters in LLMs results in a noticeable decrease in inference speed relative to real-time planners, presenting substantial obstacles to their real-world implementation.
@@ 2.1 Motion Planning For Autonomous Driving
The classic modular pipeline for autonomous driving includes perception, prediction, and planning. In this framework, the planning stage predict a future trajectory based on the perception outputs, then executed by the control system. This architecture, widely adopted in industry frameworks like Apollo [3], contrasts with end-to-end approaches [17, 16] by enabling focused research on individual tasks through well-defined data interfaces between modules.
Autonomous driving planners can be mainly categorized into rule-based and learning-based types. Rule-based planners [39, 21, 10, 11, 40, 23, 2] rely on predefined rules for determining the vehicle’s trajectory, such as maintaining a safe following distance and obeying traffic signals. For instance, IDM [39] ensures a safe distance from the leading vehicle by calculating an appropriate speed based on braking and stopping distances. PDM [10] builds on IDM by selecting the highest-scoring IDM proposal as the final trajectory, achieving state-of-the-art performance in the nuPlan Challenge 2023 [4]. However, rule-based planners often struggle with complex driving scenarios beyond their predefined rules.
Learning-based planners [20, 24, 13, 32, 33, 18] aim to replicate human expert driving trajectories using imitation learning or offline reinforcement learning from large-scale datasets. However, they face limitations due to the scope of the datasets and model complexity, leaving substantial room for improvement in areas like routing information comprehension and environmental awareness.
@@ 2.2 LLM For Autonomous Driving
The rapid advancement of Large Language Models has been noteworthy. These models, including GPT-4 [1] and Llama2 [38], have been trained on extensive textual datasets and exhibit exceptional generalization and reasoning abilities. A growing body of research has explored the application of LLMs’ decision-making capacities to the domain of autonomous driving planning [37, 7, 28, 34, 19, 44, 29, 42, 35, 36, 14, 45, 12, 8, 41, 26, 31, 27, 5, 25, 47, 43].
Some efforts [12, 29, 44, 28] have explored integrating scene information, including the ego vehicle’s status and information about obstacles, pedestrians, and other vehicles into LLMs using linguistic modalities for decision-making and explanations. These approaches face limitations due to finite context length, making it challenging to encode precise information for effective decision-making and reasoning. To overcome these constraints, multi-modal strategies such as DrivingWithLLM [7], DriveGPT4 [45], and RAGDriver [47] have been developed. These methods align vectorized or image/video modalities with linguistic instructions for a more comprehensive interpretation of driving scenarios. However, using language to express control signals has its limitations. DrivingWithLLM outputs high-level commands in linguistic expressions, improving QA interaction but reducing the fidelity of translating complex reasoning into precise vehicle control. DriveGPT4 expresses waypoints through language, showing strong open-loop performance but lacking closed-loop simulation evaluation.
Furthermore, some efforts [34, 35] focus on closed-loop evaluation by connecting low-level controllers or regressors behind large language models for precise vehicle control. LMDrive [35] uses continuous image frames and navigation instructions for closed-loop driving but requires complete LLM inference at each planning step. LanguageMPC [34] employs LLMs to obtain Model Predictive Control parameters, achieving control without training. However, these methods necessitate serial language decoding or full LLM inference at each planning step, challenging real-time responsiveness and limiting practical deployment.
In our approach, we shift from using LLMs for direct language output to enhancing real-time learning-based planners. This strategy improves environmental comprehension and allows LLMs and real-time planners to operate independently at different inference rates. This decoupling reduces LLM inference latency, facilitating real-world deployment.
@@ 3 Data Generation
The nuPlan [4] dataset is the first large-scale benchmark for autonomous driving planning, comprising $1,200$ hours of real-world human driving data from Boston, Pittsburgh, Las Vegas, and Singapore. To support various training stages, we developed pre-training and fine-tuning datasets from the nuPlan Train and Val sets, focusing on 14 official [30] challenging scenario types.
@@ 3.1 Pre-training Data Generation
To enhance LLM’s understanding of instructions in autonomous driving, we created a dataset of language-based QAs, aligning with LLM’s native modality to better grasp instruction semantics, which includes Planning-QA and Reasoning1K, with sample datasets provided in the supplementary material.
@@ 3.1.1 Planning-QA
is created using a rule-based approach for scalability. It is designed to enhance the LLM’s understanding of the relationships among waypoints, high-level instructions, and control. In this context, waypoints are arrays of points, high-level instructions are composed of velocity commands (stop, accelerate, decelerate, maintain speed) and routing commands (turn left, turn right, go straight), and control involves velocity and acceleration values. Planning-QA includes six types of questions, each focusing on the conversion between waypoints, high-level instructions, and control.
@@ 3.2 Fine-tuning Data Generation
To further achieve multimodal understanding and alignment, we constructed a fine-tuning dataset based on $10,000$ scenarios, capturing one frame every 8 seconds, resulted in a training set of $180,000$ frames and a validation set of $20,000$ frames, each incorporating both vectorized map data and linguistic prompts. Importantly, the scenario type distribution in both training and validation datasets matches the distribution of the entire nuPlan trainval dataset.
For the extracted vectorized scene information, similar to [18], ego information and $20$ surrounding agent information over $20$ historical frames, in addition to global map data centered around the ego are involved.
LLM’s prompt is comprised of two parts: system prompt and series of routing instructions. Concerning routing instructions, a rule-based approach is employed to transform the pathway into a series of instructions enhanced with distance information. Regarding training dataset preparation, the ego vehicle’s ground truth trajectory over the ensuing $8$ seconds is harnessed as the pathway for routing instruction generation. During simulation, based on the observation of the current scene, a pathway that adheres to a specified maximum length is found through a hand-crafted method as a reference path for instruction generation.
@@ 4 Methodology
As illustrated in Fig. 2, we introduce the asynchronous LLM-enhanced closed-loop framework, AsyncDriver, which mainly includes two components: 1) The Scene-Associated Instruction Feature Extraction Module; 2) The Adaptive Injection Block. Additionally, due to our framework’s design, the inference frequency between the LLM and the real-time planner could be decoupled and regulating by asynchronous interval, which markedly enhancing inference efficiency.
@@ 4.1.1 Multi-modal Input
In each planning iteration, the vectorized scene information is acquired from the simulation environment. Analogous to the approach employed by GameFormer [18], historical trajectory and state information of both the ego and other agents are extracted, alongside global map data. Vectorized scene information for the real-time planner is provided in the same manner. All vector data are relative to the position of the ego. Subsequently, through the processing by the Vector Map Encoder and Map Adapter, we derive map embeddings. These map embeddings, along with language embeddings, are then fed into the Llama2-13B backbone to obtain the final hidden features $h=\{h_{0},h_{1},...,h_{-1}\}\in\mathbb{R}^{N_{h}\times D_{llm}}$.
@@ 4.2 Adaptive Injection Block
We adopt the decoder structure of [18] as our basic decoder layer, facilitating the adaptive integration of scene-associated instruction features by evolving the conventional transformer-based decoder layer into an Adaptive Injection Block.
@@ 4.4 Training Details
Following [18], the planning loss comprises two parts: 1) Mode prediction loss. $m$ different modes of trajectories of neighbor agents are represented by Gaussian mixture model (GMM). For each mode, at any given timestamp $t$, its characteristics are delineated by a mean ${\mu}^{t}$ and covariance ${{\sigma}^{t}}$ forming a Gaussian distribution. The best mode ${m}^{*}$ is identified through alignment with the ground truth and refined via the minimization of negative log-likelihood. 2) Ego Trajectory Prediction loss. Future trajectory points of ego vehicle are predicted and are refined by $l1$ loss. Consequently, the planning loss is as follows:
@@ 5.1.1 Evaluation Settings
In accordance with the nuPlan challenge $2023$ [30] settings, we selected $14$ official challenging scenario types for training and evaluation. While nuPlan [4] includes $757,844$ scenarios, most simple scenarios are insufficient for critical planner performance assessment, and the vast data volume extends evaluation times. We chose the Hard20 dataset by randomly selecting $100$ scenarios per type from the test set, then using the PDM [10] planner (nuPlan $2023$ champion) to retain the $20$ lowest-scoring scenarios per type, resulting in a test set of $279$ scenarios.
@@
UrbanDriver [33] 35.35 75.53 97.12 98.56 85.23 55.21 81.62 47.84
GCPGP [13] 36.85 81.29 98.20 77.33 46.96 72.30 97.92 68.34
IDM [39] 53.07 84.94 98.02 83.15 64.79 74.01 96.38 60.57
GameFormer [18] 62.05 93.54 98.74 83.15 66.27 86.02 98.19 74.55
PDM-Hybird [10] 64.05 95.34 99.10 75.98 67.93 87.81 99.57 72.75
PDM-Closed [10] 64.18 95.69 99.10 77.06 68.20 87.81 99.57 73.47
@@
Table 2: Scores per scenario types in evaluation on nuPlan Closed-Loop Reactive Challenges on Hard20 split. The best results are highlighted in bold , while the second-best results are underscored with an underline for clear distinction. Types 00-$13$ represent the $14$ official scenario types of the nuPlan challenge 2023 [30], with specific details provided in the supplementary materials.
@@
UrbanDriver [33] 69.39 15.78 44.59 7.42 13.7 27.14 0.00 19.44 20.8 23.61 68.33 93.16 33.78 56.39
GCPGP [13] 59.50 35.91 33.6 29.33 42.84 17.24 4.86 32.33 8.22 36.23 71.32 80.46 36.19 23.61
IDM [39] 70.69 44.84 91.54 54.08 50.22 41.71 4.76 53.97 37.53 60.33 83.97 93.03 45.77 12.44
GameFormer [18] 84.32 65.78 83.62 49.03 71.79 36.8 0.00 51.76 55.03 77.24 82.83 98.24 49.41 56.16
PDM-Hybird [10] 87.68 69.61 87.20 49.95 82.80 41.32 4.23 54.98 51.72 82.95 80.37 98.40 37.77 71.99
PDM-Closed [10] 87.67 70.93 87.20 49.95 82.80 41.25 4.22 53.53 51.57 82.96 80.37 98.40 39.95 72.04